# ZeroTuning: Unlocking the Initial Token’s Power to Enhance Large Language Models Without Training

### v1

@@ Abstract
Training-free methods for enhancing large language models (LLMs) have attracted growing interest recently, with token-level attention tuning emerging as an interpretable and promising direction. However, existing methods typically rely on auxiliary mechanisms to identify important or irrelevant task-specific tokens, introducing potential bias and limiting applicability. In this work, we uncover a surprising and elegant alternative: the semantically empty initial token (e.g., <BOS> in Llama) serves as a powerful and underexplored control point for optimizing model behavior. Through theoretical analysis, we show that tuning the initial token’s attention sharpens or flattens the attention distribution over subsequent tokens, and its role as an attention sink amplifies this effect. Empirically, we find that: (1) tuning its attention improves LLM performance across tasks more effectively than tuning other task-specific tokens; (2) the effect follows a consistent trend across layers, with earlier layers having greater impact, but varies across attention heads, with different heads showing distinct preferences in how they attend to this token. Based on these findings, we propose ZeroTuning, a training-free approach that improves LLM performance by applying head-specific attention adjustments to this special token. Despite tuning only one token, ZeroTuning achieves higher average performance on text classification, multiple-choice QA, and multi-turn conversation tasks across models such as LLama, Qwen, and DeepSeek. For example, ZeroTuning improves Llama-3.1-8B by 11.71% on classification tasks, 2.64% on QA tasks, and raises its multi-turn score from 7.804 to 7.966. The method is also robust to limited resources, few-shot settings, long contexts, quantization, decoding strategies, and prompt variations. Our work sheds light on a previously overlooked control point in LLMs, offering new insights into both inference-time tuning and model interpretability.
@@ 1 Introduction
Recent research has explored the potential of lightweight and training-free methods to adapt pre-trained large language models (LLMs) for downstream applications. One promising direction is token-level attention tuning, which modifies the attention distribution at inference time without changing any parameters. Unlike fine-tuning [1, 2, 3, 4, 5, 6] or prompt engineering [7, 8, 9], which treat LLMs as a black box, token-level attention tuning introduces a more transparent way to guide the model’s focus by increasing attention to key tokens or downplaying irrelevant ones. These methods can even outperform prompt-based approaches in certain scenarios, such as open-domain question answering [10]. Examples include PASTA (manual token selection) [11], AutoPASTA (LLM-guided token identification) [10], and ACT (attention-thresholding-based sink token identification) [12]. Similar ideas have also been applied to vision-language models to mitigate hallucinations by increasing attention to image tokens [13, 14, 15].
While effective, these methods require external mechanisms to identify task-specific tokens, which may introduce potential bias (e.g., emphasizing misleading tokens) and limit applicability when the importance of tokens is unclear or attention maps are not accessible. This motivates a question: Is it possible to improve model performance by tuning the attention to a universal and task-agnostic token without relying on task-specific token identification?
In this paper, we reveal that the answer lies in a widely existing yet often overlooked token: the initial token (e.g., <BOS> in LLaMA). Though it is commonly researched and regarded as an attention sink [16, 17, 18, 19], prior tuning methods have largely ignored it, focusing instead on visible input tokens. Our theoretical analysis shows that tuning the attention of this special token can sharpen or flatten the attention distribution across subsequent task-relevant tokens, while preserving their relative importance. Its role as an attention sink further amplifies this effect (Section 2.1). Additionally, we demonstrate that modifying the initial token’s key or query states achieves similar effects when direct tuning of attention weights is infeasible.
We conducted a series of experiments to investigate the impact of tuning the initial token’s attention and found the following:
@@
    Despite lacking semantic meaning, the initial token acts as an effective control point for steering model behavior: tuning its attention consistently improves LLM performance more effectively than tuning any other token (Section 2.2).
@@
    The tuning effect propagates consistently across layers, with the shallow and middle layers contributing more than the deeper ones; jointly tuning all layers yields the best results (Section 2.3).
@@
    The impact also varies across attention heads: some heads respond positively to increased attention on this token, while others respond negatively. We categorize these as up-effective and down-effective heads, and show that selectively tuning them outperforms uniform tuning across all heads (Sections 2.4 and 2.5).
These findings motivate ZeroTuning, a simple and effective training-free method that recalibrates the initial token’s attention to improve the LLM’s performance without requiring task-specific token identification (Section  2.6). Experiments demonstrate that, despite tuning only one token, ZeroTuning achieves higher average performance on text classification (e.g., SST2, BoolQ), multiple-choice QA (e.g., MMLU, LogiQA), and multi-turn conversation (e.g., MT-Bench) benchmarks across LLMs including Llama-3.1-8B-Instruct, Llama-2-13B-Instruct, Qwen-2-7B, and Deepseek-R1-14B. For example, ZeroTuning improves Llama-3.1-8B by 11.71% on classification tasks, 2.64% on QA tasks, and raises its multi-turn score from 7.804 to 7.966. The method is also robust under various conditions, including limited resources, few-shot settings, long contexts, quantization, decoding strategies, and prompt variations.
Our work sheds light on a previously overlooked control point in LLMs, offering new insights into both inference-time tuning and model interpretability.
@@ 2 Unveiling the Power of the Initial Token
In this section, we delve into the impact of tuning the initial token’s attention and introduce our method. Unless otherwise specified, all experiments use the Llama-3.1-8B-Instruct model, with dataset details provided in Section 3.1.
@@ 2.1 Formalizing the Tuning Process
Figure 1: Effect of the scaling factor $\gamma\in[0,10]$ and initial token’s attention $a_{0}\in\{0,0.2,0.4,0.6,0.8\}$ on attention differences $|a_{i}^{\prime}-a_{j}^{\prime}|$ between non-initial tokens. Shaded regions highlight tuning regimes: excessive tuning ($\gamma\in[0,0.3]\cup[6.5,10]$) that may impair LLM performance, moderate sharpening ($\gamma\in[0.3,1]$) that amplifies token-level attention differences, and moderate smoothing ($\gamma\in[1,6.5]$) that reduces them. Decreasing $\gamma$ sharpens the distribution by amplifying attention differences, while increasing $\gamma$ smooths it by reducing disparities.
In decoder-only Transformers, the next token is generated using the final token as the query, attending to all preceding tokens (including itself) as keys. Consider a single-layer, single-head attention mechanism with an input sequence:
$\mathbf{X}=[x_{0},x_{1},\dots,x_{T-1}]\in\mathbb{R}^{d\times T}.$ (1)
At generation step $T$, the query is $x_{T-1}$, and the attention weights over the keys $x_{0}$ to $x_{T-1}$ form a probability vector:
$\boldsymbol{a}=[a_{0},a_{1},\dots,a_{T-1}],\quad a_{i}\geq 0,\quad\sum_{i=0}^{T-1}a_{i}=1,$ (2)
where $a_{0}$ denotes the attention of the initial token, while $a_{1}$ to $a_{T-1}$ correspond to subsequent tokens.
To control the influence of $x_{0}$, we scale its attention weight by a factor $\gamma>0$ and re-normalize:
@@
where the normalization constant is
$D=\gamma a_{0}+\sum_{i=1}^{T-1}a_{i}=\gamma a_{0}+(1-a_{0})=(\gamma-1)a_{0}+1.$ (4)
This rescaling preserves the relative proportions among non-initial tokens:
@@
Intuitively, setting $\gamma>1$ amplifies $a_{0}$ and flattens the attention distribution over the remaining tokens. Conversely, $\gamma<1$ suppresses $a_{0}$ and sharpens distinctions among other tokens. When $\gamma=1$, the attention remains unchanged. As illustrated in Figure 1, the value of $a_{0}$ determines the strength of the tuning effect—larger $a_{0}$ values lead to more pronounced modulation. However, extreme values of $\gamma$ can be detrimental: overly small values degrade model performance, a behavior consistent with prior findings on attention sink studies [19], while excessively large values can also hurt performance by suppressing attention to downstream tokens toward near-zero values.
Then, we analyze how this tuning influences the final representations. Let the value states corresponding to each token be:
$\mathbf{V}=[\mathbf{v}_{0},\mathbf{v}_{1},\dots,\mathbf{v}_{T-1}]\in\mathbb{R}^{d_{v}\times T},$ (7)
where $\mathbf{v}_{i}\in\mathbb{R}^{d_{v}}$ is the value vector associated with token $x_{i}$. The attention output is a weighted sum of these value vectors. Applying the scaled attention weights yields:
$\mathbf{o}$ $=\sum_{i=0}^{T-1}a_{i}^{\prime}\mathbf{v}_{i}=\frac{\gamma a_{0}}{D}\mathbf{v}_{0}+\sum_{i=1}^{T-1}\frac{a_{i}}{D}\mathbf{v}_{i},$
$=\frac{1}{D}\left[\gamma a_{0}\mathbf{v}_{0}+\sum_{i=1}^{T-1}a_{i}\mathbf{v}_{i}\right].$ (8)
Previous studies have observed the “value-state drain” phenomenon, wherein the value state of the initial token tends to be much smaller than those of the other tokens on average [20, 21, 22, 18]. Therefore, the contribution of the first term $\gamma a_{0}\mathbf{v}_{0}$ to the final output is relatively small, and the value states of the subsequent tokens dominate the final representation. Tuning $\gamma$ adjusts the relative contributions of these value states, influencing the final representation.
@@ 2.2 The Unique Importance of the Initial Token
Given the special role of the initial token in LLMs, we begin by addressing two key questions: (a) Can tuning its attention improve performance across downstream tasks? (b) Is this position more effective and influential than others? To investigate, we conduct a controlled experiment in which we uniformly scale the attention scores of a single token position across all heads and layers using a scaling factor $\gamma$. We evaluate the resulting performance on three downstream tasks: SST-2, BoolQ, and LogiQA. For comparison, we repeat the same procedure for other positions, including the second, third, middle ($\lfloor T/2\rfloor$), and final tokens. As shown in Figure 2, tuning the attention of the initial token consistently yields the largest and most stable performance gains across all tasks. Interestingly, performance varies with the direction of tuning: tasks like SST-2 benefit from up-scaling ($\gamma>1$), while BoolQ and LogiQA improve with down-scaling ($\gamma<1$).
Previous work has identified the initial token as an attention sink that helps prevent over-mixing of information during autoregressive generation [18, 19]. We extend this understanding by showing that tuning the attention of the initial token can reshape the attention distribution and improve model performance across tasks. For example (see Appendix L), in the SST-2 sentiment classification task, input sentences often contain both positive and negative cues. We find that LLMs tend to overemphasize isolated negative keywords, leading to biased predictions. By increasing attention on the initial token, we flatten the attention distribution, promoting a more balanced integration of information and fostering a more holistic understanding. In contrast, in the BoolQ task, where long passages must be inspected for a small number of critical cues, the model may miss relevant tokens due to excessively diffuse attention. Reducing initial tokens’ attention amplifies the relative differences between the remaining tokens, sharpening focus on salient input regions and improving response accuracy.
As discussed in Section 2.1, the higher the attention on the initial token, the greater the impact of such tuning. Given that attention sinks are typically concentrated at this position, the initial token is inherently more impactful when modulated, explaining its outsized influence relative to other tokens. Besides, unlike other tokens, the initial token is task-agnostic; adjusting its attention does not alter the relative importance of task-relevant tokens, thereby avoiding harmful biases such as inadvertently amplifying attention to irrelevant content.
@@ 2.3 Layer-wise Analysis of Initial Token Scaling
To better understand the impact of tuning the initial token’s attention, we examine how its effect varies when applied selectively across different layers. Following prior work on layer functionality in transformer-based models [23, 24], we divide the 32 layers of Llama-3.1-8B-Instruct into three groups: shallow layers (Layers 1–10), middle layers (Layers 11–21), and deep layers (Layers 22–31). We then perform independent attention scaling experiments for each group on six tasks: BoolQ, SST-2, SST-5, MR, LogiQA, and MathQA. Based on earlier findings, we apply a scaling range of $[0,1]$ for BoolQ and LogiQA, and $[1,2]$ for the remaining tasks.
@@
As shown in Figure 3, tuning the initial token’s attention yields consistent trends across all three layer groups (i.e., accuracy changes similarly with scaling regardless of depth), and jointly tuning all layers amplifies these benefits, often resulting in the highest accuracy. However, the magnitude of improvement varies. In most cases, tuning the shallow and middle layers leads to greater accuracy than tuning the deep layers. This aligns with prior findings that early and middle layers primarily contribute to representation learning and knowledge integration, whereas deep layers focus on task-specific reasoning over aggregated features [25, 26].

### v2

@@ Abstract
Token-level attention tuning—a class of training-free methods including Post-hoc Attention Steering (PASTA) and Attention Calibration (ACT)—has emerged as a promising way to improve frozen LLMs with interpretable interventions. However, these methods depend on auxiliary heuristics to identify “important” task-specific tokens, which can introduce bias and limit applicability when token importance is unclear or when using optimized kernels where attention maps are inaccessible. We propose a simpler and more elegant alternative: acting only on the initial token (e.g., <BOS> in LLaMA). We show theoretically that adding lightweight biases to this token’s attention logits monotonically controls the entropy of the downstream attention distribution—an effect amplified by its natural function as an attention sink. Our empirical analysis reveals that this tuning process can positively affect LLMs and better unlock their pretrained knowledge, with stronger effects in early layers and distinct scaling preferences across attention heads. Building on these insights, we introduce ZeroTuning: a training-free method that improves LLM performance by applying head-specific attention adjustments to the initial token, requiring zero parameter updates. We present two variants: a supervised mode that calibrates on validation examples, and a novel unsupervised mode that directly minimizes the model’s output entropy. Our method requires no KV‑cache or decoding changes, and is kernel‑agnostic (works with SDPA and FlashAttention). The method is lightweight and requires only four lines of modification to standard LlamaAttention code. It achieves broad gains across 15 datasets and outperforms previous, more complex methods; for instance, with Llama-3.1-8B, it yields relative improvements of 19.9% on classification, 4.5% on question answering, and 2.1% on dialogue. ZeroTuning also works out-of-the-box with quantized inference and maintains its performance improvements with increasing context lengths. Our code and runnable demo are available at https://anonymous.4open.science/r/ZeroTuning.
@@ 1 Introduction
Training-free methods have been widely explored to enhance Large Language Models (LLMs) performance at inference time. Among these, token-level attention tuning has emerged as a particularly promising direction, offering an interpretable way to steer model behavior by modifying the attention distribution without any parameter updates. Unlike fine-tuning (Hu et al., 2021; Dettmers et al., 2023) or prompt engineering (Wei et al., 2022; Wang et al., 2022; Han et al., 2025a; b), which largely treat LLMs as black boxes, attention tuning provides a transparent mechanism to guide the model’s focus. Methods such as Post-hoc Attention Steering (PASTA (Zhang et al., 2023a), AutoPASTA (Zhang et al., 2024a)), and Attention Calibration (ACT) (Yu et al., 2024) have demonstrated the power of this approach, even outperforming prompting-based techniques in complex tasks like open-domain question answering (Zhang et al., 2024a). Similar principles have been successfully applied to vision-language models to mitigate hallucinations by re-weighting attention towards image tokens (Liu et al., 2024b; Zhu et al., 2024; Wei & Zhang, 2024).
However, the efficacy of these methods is fundamentally constrained by their reliance on external, often heuristic, mechanisms to identify task-specific ”important” tokens. This dependency not only introduces the risk of bias (e.g., amplifying misleading cues) but also limits their applicability in scenarios where token importance is ambiguous or when using optimized attention kernels where attention maps are inaccessible. This critical limitation motivates a fundamental question: Is it possible to enhance model performance by tuning a universal, task-agnostic token, thereby bypassing the need for fragile, task-specific token identification?
Figure 1: Overview of the ZeroTuning method and its effectiveness. Left: Our method requires only a few lines of code to scale the initial token’s attention within the model’s forward pass. Right: This simple intervention yields significant and consistent improvements across a variety of LLMs and tasks.
In this paper, we demonstrate that the answer is a definitive yes. The solution lies not in adding complexity, but in leveraging a ubiquitous yet often overlooked architectural artifact: the initial token (e.g., <BOS> in LLaMA). While its tendency to function as an ”attention sink” is well-documented (Xiao et al., 2023; Kaul et al., 2024; Gu et al., 2024; Barbero et al., 2025), its potential as an active control lever for performance enhancement has remained largely untapped.
Our theoretical analysis reveals that modulating the attention on this single token allows for precise control – sharpening or smoothing – over the entire downstream attention distribution while preserving the relative importance of task-relevant tokens. Its natural role as an attention sink further amplifies this effect (Sec 3.1). Furthermore, we show this control can be achieved even without direct access to attention weights by modifying the initial token’s key or query states.
To validate this insight, we conducted a systematic investigation and uncovered three key findings:
@@
    Despite lacking semantic meaning, the initial token acts as a positive and effective control point for steering model behavior. Tuning its attention corrects the model’s biases and unlocks its pretrained knowledge with reduced output entropy. This approach consistently yields greater performance gains than tuning any other token (Sec 3.2).
@@
    The influence of this tuning is systemic, propagating consistently across layers. The shallow and middle layers are the most impactful, though jointly tuning all layers produces the strongest results (Sec 3.3).
@@
    The effect is heterogeneous across attention heads; some respond positively to increased attention (up-effective), while others respond negatively (down-effective). We show that selectively targeting the dominant head type outperforms uniform tuning (Sec 3.4,Sec 3.5).
Building on these findings, we introduce ZeroTuning (see Figure 1), a simple, powerful, and training-free method that recalibrates the initial token’s attention to boost LLM performance without any task-specific identification (Sec 3.6). We introduce two variants for attention calibration: a supervised mode by maximizing the accuracy on the labeled validation set, and a novel unsupervised mode by minimizing output entropy. Across a suite of 15 benchmarks, ZeroTuning achieves substantial gains on models like Llama-3.1-8B-Instruct, Llama-2-13B-Instruct, Qwen-2-7B, and Deepseek-R1-14B. For instance, it boosts Llama-3.1-8B-Instruct performance by a relative 19.9% on classification, 4.5% on question answering, and raises its MT-Bench score from 7.804 to 7.966. The method demonstrates remarkable robustness across long contexts, few-shot settings, quantization, and prompt variations. Our work not only delivers a practical tool for lightweight model enhancement but also sheds new light on a fundamental control mechanism within LLMs, advancing both inference-time optimization and model interpretability.
2 Related Work
Our work is situated at the intersection of two active research areas: inference-time attention tuning and the mechanistic understanding of initial tokens. A growing body of work has shown that modifying token-level attention at inference time can enhance the performance of both LLMs and VLMs (Yu et al., 2024; Zhang et al., 2023a; Liu et al., 2024b; Wei & Zhang, 2024). However, prevailing methods like PASTA (Zhang et al., 2023a) and Auto-PASTA (Zhang et al., 2024a), which identify and up-weight key tokens, or ACT (Yu et al., 2024), which down-weights non-initial sink tokens, fundamentally rely on heuristics to identify task-specific tokens. This reliance limits their universality and introduces potential biases. Concurrently, another line of research has focused on explaining why the initial token often becomes an ”attention sink” (Xiao et al., 2023), attributing it to architectural biases and its role as a stabilizing anchor (Barbero et al., 2025; Gu et al., 2024). While these studies provide a crucial understanding of what the phenomenon is, the question of how to actively and elegantly harness it for performance gains remains largely unexplored. Our work bridges this gap. We shift the focus from task-specific token identification to a universal, task-agnostic control point, and move from passive observation of the initial token to a practical tuning framework that leverages its unique properties. Detailed related work is provided in Appendix A.
@@ 3 Unveiling the Power of the Initial Token
In this section, we first formalize the mechanism of tuning the initial token’s attention, then empirically demonstrate its unique importance and dissect its effects across the model’s layers and heads. This systematic analysis culminates in our proposed ZeroTuning methodology. Unless otherwise specified, all experiments use the Llama-3.1-8B-Instruct model, with setup details in Section 4.1.
@@ 3.1 Formalizing the Tuning Process
In a decoder-only Transformer, autoregressive generation for a sequence $\mathbf{X}=[x_{0},x_{1},\dots,x_{T-1}]\in\mathbb{R}^{d\times T}$ involves a causal self-attention mechanism. At timestep $T$, the query is derived from the final token representation, $x_{T-1}$, which attends to all preceding token representations (including itself) as keys. This process yields an attention weight distribution over the input sequence:
$\bm{a}=[a_{0},a_{1},\dots,a_{T-1}],\quad\text{where}\quad a_{i}\geq 0\quad\text{and}\quad\sum_{i=0}^{T-1}a_{i}=1.$ (1)
Here, $a_{0}$ is the attention score assigned to the initial token, while $a_{1},\dots,a_{T-1}$ correspond to subsequent tokens. To control the influence of $x_{0}$, we introduce a tuning factor $\gamma>0$ to scale its attention and re-normalize:
@@
where the normalization constant $D=\gamma a_{0}+\sum_{i=1}^{T-1}a_{i}=(\gamma-1)a_{0}+1$.
This rescaling preserves the relative proportions among all non-initial tokens:
@@
Intuitively, $\gamma>1$ amplifies $a_{0}$, flattening the remaining distribution, while $\gamma<1$ suppresses $a_{0}$, sharpening it.
Theoretically, the magnitude of this effect is governed by the initial token’s own attention weight, $a_{0}$. We define this effect, $E_{\text{diff},i,j}$, as the change in attention difference between any two non-initial tokens $i$ and $j$:
$E_{\text{diff},i,j}=|(a_{i}^{\prime}-a_{j}^{\prime})-(a_{i}-a_{j})|=|a_{i}-a_{j}|\left|\frac{1}{(\gamma-1)a_{0}+1}-1\right|=|a_{i}-a_{j}|\frac{|\gamma-1|a_{0}}{(\gamma-1)a_{0}+1}.$ (5)
To analyze how $E_{\text{diff},i,j}$ varies with $a_{0}$, we take its partial derivative with respect to $a_{0}$:
$\frac{\partial E_{\text{diff},i,j}}{\partial a_{0}}=|a_{i}-a_{j}||\gamma-1|\cdot\frac{1}{((\gamma-1)a_{0}+1)^{2}}.$ (6)
Given that $|a_{i}-a_{j}||\gamma-1|\geq 0$ and the denominator $((\gamma-1)a_{0}+1)^{2}=D^{2}>0$, the derivative is non-negative. Thus, $E_{\text{diff},i,j}$ is a monotonically non-decreasing function of $a_{0}$. In any non-trivial case ($\gamma\neq 1$ and $a_{i}\neq a_{j}$), the relationship is strictly increasing. A detailed proof and visualization are in Appendix C.
This result provides a crucial insight: the larger the initial token’s attention score ($a_{0}$), the more powerful it becomes as a lever for controlling the entire attention distribution. Since prior work has established that initial tokens are natural attention sinks (Barbero et al., 2025), they are inherently potent control points for this tuning process.
@@ 3.2 The Unique Importance of the Initial Token
Given the special role of the initial token, we first investigate a key empirical question: (a) Does tuning its attention positively impact performance on downstream tasks? (b) Is this position more effective and influential than others? To investigate, we conduct a controlled experiment in which we uniformly scale the attention scores of a single token position across all heads and layers using a scaling factor $\gamma$. We evaluate the resulting performance on three downstream tasks: SST-2, BoolQ, and LogiQA. For comparison, we repeat the same procedure for other positions, including the second, third, middle ($\lfloor T/2\rfloor$), and final tokens. As shown in Figure 2, tuning the attention of the initial token consistently yields the largest and most stable performance gains across all tasks. Interestingly, performance varies with the direction of tuning: tasks like SST-2 benefit from up-scaling ($\gamma>1$), while BoolQ and LogiQA improve with down-scaling ($\gamma<1$).
Figure 3: The impact of the tuning.
Previous work has identified the initial token as an attention sink that helps prevent over-mixing of information during autoregressive generation (Gu et al., 2024; Barbero et al., 2025). Our empirical results extend this understanding, showing that tuning the initial token’s attention beneficially reshapes the distribution over subsequent tokens and, in turn, improves model performance across tasks. This beneficial effect can be understood from two perspectives.
1. Correcting Biases. The tuning process counteracts reasoning flaws caused by pretrained attention biases for each dataset, as visualized in Figure 3. Our error analysis reveals a clear pattern: (a) Up-scaling ($\gamma>1$) benefits tasks requiring holistic context integration. It flattens the attention distribution, preventing the model from over-focusing on misleading local cues. For instance, in SST-2 sentiment analysis (see Appendix F), LLMs often fixate on isolated negative keywords while ignoring the surrounding positive context. Increasing the initial token’s attention promotes a more balanced, global understanding, thereby correcting such biased predictions. (b) Down-scaling ($\gamma<1$) benefits tasks demanding sharp focus on critical evidence. It amplifies the relative weights of non-initial tokens, helping the model pinpoint key details in a diffuse context. In long-context tasks like BoolQ, a model’s focus can become diluted, failing to locate the precise text segment containing the answer. Reducing the initial token’s attention sharpens the model’s focus on this salient information.
Figure 4: A strong inverse correlation between accuracy (orange) and next-token prediction entropy (blue).
2. Reducing Predictive Uncertainty. The tuning process can be viewed through the lens of output entropy, a proxy for model uncertainty. As illustrated in Figure 4, a clear inverse correlation emerges: the scaling factor that minimizes entropy consistently aligns with the factor that maximizes accuracy. This suggests our method better unlocks the model’s pretrained knowledge, leading to more confident and correct predictions.
@@ 3.3 Layer-wise Analysis of Initial Token Scaling
To understand the propagation of this effect, we examine how its effect varies when applied selectively across different layers. Following prior work on layer functionality in transformer-based models (Jin et al., 2024; Zhang et al., 2024b), we divide the 32 layers of Llama-3.1-8B-Instruct into three groups: shallow (Layers 1–10), middle (Layers 11–21), and deep (Layers 22–31). We then perform independent attention scaling experiments for each group on six tasks: BoolQ, SST-2, SST-5, MR, LogiQA, and MathQA. Based on earlier findings, we apply a scaling range of $[0,1]$ for BoolQ and LogiQA, and $[1,2]$ for the remaining tasks.
@@
As shown in Figure 5, tuning yields consistent trends at any depth (i.e., accuracy changes similarly with scaling regardless of depth), and jointly tuning all layers amplifies these benefits, often resulting in the highest accuracy. However, the magnitude of improvement varies. In most cases, tuning the shallow and middle layers leads to greater accuracy than tuning the deep layers.
Prior studies have found that early and middle layers mainly support representation learning and knowledge integration, while deep layers focus on task-specific reasoning over aggregated features (Chen et al., 2024; Jin et al., 2025). Therefore, we argue that the tuning process more effectively reshapes the representational space in shallow and middle layers, promoting better downstream performance and reducing uncertainty.